# ExGRPO: Learning to Reason from Experience

### v1

@@ 1 Introduction
Experience replay (Lin, 1992) is a widely adopted technique in RL to address this issue and improve sample efficiency. As stated in Silver & Sutton (2025), AI is at the cusp of a new period in which experience will become the dominant medium of improvement. This idea, often referred to as experience-based RL, leverages the intuition that previously collected interactions (i.e., state-action-reward tuples) contain valuable information for learning, helping models stabilize training and accelerate convergence (Mnih et al., 2013). A non-trivial challenge lies in how to exploit past experiences based on their differing “values” (Schaul et al., 2016) and manage the replay process according to customized learning schedules (Sujit et al., 2023). However, efficient experience replay mechanisms remain largely underexplored in RLVR for building large reasoning models (LRMs;  Plaat et al. 2024; Zhang et al. 2025b; Qu et al. 2025). Given the vast quantity of experiences collected during rollouts, a fundamental question remains: How can the reasoning model effectively exploit its own stream of experience to maximize learning toward scaling RL compute for LRMs?
@@ 2 Related Work
Reinforcement Learning with Verifiable Rewards (RLVR; Lambert et al. 2024) frames the language model as a policy that generates reasoning trajectories for verifiable tasks, with rewards provided by rule-based (Guo et al., 2025; Yu et al., 2025) or model-based verifiers (Su et al., 2025; Ma et al., 2025b; Chen et al., 2025). Most RLVR methods adopt on-policy optimization (Schulman et al., 2017; Shao et al., 2024), which ensures stability but incurs high computational cost. Recent work has explored off-policy techniques (Kallus & Uehara, 2020; Meng et al., 2023) that incorporate historical or external data. Some approaches mix expert demonstrations with on-policy updates, e.g., off-policy policy gradients (Yan et al., 2025), SFT loss (Zhang et al., 2025c; Ma et al., 2025a), or knowledge distillation (Xu et al., 2025); others develop direct off-policy update rules for improved sample efficiency (Cohen et al., 2025; Roux et al., 2025; Arnal et al., 2025). However, they overlook the impact of data quality within the experience replay buffer, a factor that remains underexplored in RLVR and is the central focus of our work.
@@
Experience replay (Lin, 1992) is a classical RL technique to improve sample efficiency and stabilize training, later extended with prioritized replay (Schaul et al., 2016; Sujit et al., 2023) to emphasize informative transitions, enabling breakthroughs in control tasks (Mnih et al., 2013; 2015; Lillicrap et al., 2016; Zha et al., 2019). For LRMs, recent studies show that replaying successful trajectories accelerates convergence and improves reasoning capabilities. ReMix (Liang et al., 2025) leverages off-policy experience replay to improve training efficiency; RePO adopts online replay (Li et al., 2025), collecting early on-policy rollouts and revisiting them asynchronously; RLEP (Zhang et al., 2025a) reuses trajectories from a well-trained policy; and RRL (Dou et al., 2025) dynamically revisits promising early states; ARPO (Lu et al., 2025) extends this idea to GUI agents, combining GRPO with a replay buffer to enhance stability and sample efficiency. Among them, the last three methods overlook importance weighting to correct distribution mismatch in off-policy updates. In this work, we argue that not all stored experiences are equally valuable in zero RLVR for LRMs. We systematically analyze properties that determine the value of past experiences and design methods to better leverage high-value experiences, thereby improving efficiency and performance.
@@ 3.2 Preliminary Study on Experience Data
Figure 1: Analysis of question difficulty and entropy in on-policy RLVR training: (a) Test performance of models trained on different question groups; (b) Entropy distributions of logically correct trajectories across question groups; (c) Entropy comparison between correct and incorrect trajectories.
@@
Prior work has shown that the singular pursuit of minimal entropy in RLVR can lead to entropy collapse (Cui et al., 2025b), thus motivating the need for a nuanced exploration–exploitation trade-off in the context of experience management. While high-entropy policies encourage exploration, they can yield low-quality trajectories, many successes are merely lucky hits from incorrect reasoning chains (Wen et al., 2025).
@@ 4.1 ExGRPO: Experience Management
The replay buffer $\mathcal{E}$ is partitioned into buckets by each question’s latest correctness rate $\mathrm{Acc}(q^{*})$, with $q^{*}$ as the indexing key. This bucketing mechanism categorizes experiences by difficulty. To prevent overfitting on trivial cases, we introduce a Retired Set: questions solved in all rollouts are removed from $\mathcal{E}$, ensuring optimization focuses on partially solved questions.
@@
(2) Trajectory Selection. For each sampled question, we select the trajectory with the lowest entropy under the current policy, i.e., $o^{*}\leftarrow\arg\min_{o_{i}\in\{o^{*}\}}H(o_{i};\pi_{\theta})$, thereby prioritizing trajectories which indicate better CoT quality as discussed in Section 3.2.
@@ 4.2 ExGRPO: Experiential Policy Optimization
The on-policy objective $\mathcal{J}_{\text{on}}(\theta)$ follows GRPO in Eq. 3: for each query $q\sim\mathcal{B}_{\text{on}}$, $K$ trajectories $\{o_{i}\}_{i=1}^{K}$ are sampled from policy $\pi_{\theta_{\text{old}}}$ to form a advantage group $\mathcal{G}_{q}$, with objective $\mathcal{J}_{\text{GRPO}}(q,\mathcal{G}_{q};\theta)$.
@@ 5 Experiments
To assess the effectiveness of ExGRPO, we explore the following key questions: (1) How does ExGRPO perform under different training scenarios? (2) Can ExGRPO generalize across models with different capacities? (3) Can ExGRPO generalize beyond reasoning-oriented training data?
@@ 6 Analysis and Discussion
To understand the effectiveness of ExGRPO, we investigate the following research questions: (1) How does experience replay influence training dynamics, especially for different models? (2) What is the impact of experience selection on overcoming performance bottlenecks? (3) How do the components of ExGRPO, e.g., experience management, contribute to overall performance?
@@
Figure 5: Dynamics of experience replay buffer and retried set.
@@
The efficiency of experience utilization, rather than its sheer volume, is the primary driver of model performance. Figure 6 illustrates the dynamics of the experience buffer and the retired set under different data conditions. When trained on the same amount of data, the absence of question selection mechanisms (e.g., ExGRPO vs. w/o Ques. Sel.) leads to growth of the retired set and a corresponding reduction in buffer size (primarily due to repeated exposure to easy questions).
Figure 6: Dynamics of experience under different data conditions.
As a result, experience replay becomes inefficient, with high consumption of samples but low return in terms of performance gains. We further examine the scaling effects of experience replay by varying the experience proportion $\rho$ in batch optimization. An excessively high replay ratio ($\rho=75\%$) traps the model in experience exploitation. This stifles exploration, leading to a smaller experience buffer and a reduced retired set, and ultimately causing a worse performance even below the on-policy baseline. Conversely, at $\rho=25\%$, training is driven primarily by fresh on-policy learning, with replay serving as a complementary signal. Although this setting enlarges the buffer, the retired set grows only marginally, suggesting that additional experiences do not necessarily translate into better capabilities. These observations demonstrate that the utility of experience replay depends not on maximizing the buffer size, but on balancing exploration with efficient reuse of past samples, highlighting the necessity of principled experience management. Furthermore, experimental results in Appendix Table 4 indicate that optimal performance is achieved when a balance (i.e., $\rho=50\%$) is struck between replaying past experiences and learning from new data.
@@
We conduct an ablation study to assess the contributions of core components of ExGRPO. Figure 7 shows performance dynamics of validation set (see Section E.1). We evaluate the impact of removing Quality (Q.) and Trajectory (T.) selection strategies and policy shaping. For both question and trajectory selection, our analytically derived heuristics consistently outperform random selection, while random selection itself offers marginal gains over the on-policy baseline. Policy shaping mitigates this by ensuring exploitation does not hinder exploration. It is also noteworthy that applying shaping alone in on-policy RLVR yields no improvement (Yan et al., 2025), confirming that ExGRPO’s gains cannot be attributed to naive shaping. Besides, additional ablations and test-set results can be found in Section E.4. Replacing our selection method with an alternative heuristic, i.e., selecting the highest-entropy trajectory, results in degraded performance. Similarly, removing trajectory selection and importance-sampling correction mechanisms also leads to inferior performance, though the effect is less severe than other ablations. The above results and additional ablations on test set validate the rationale behind our heuristic design of experience management.
@@ 7 Conclusion
While experiential RL offers a promising approach to mitigate sample inefficiency in RLVR for large reasoning models, this area still lacks systematic exploration of experience value and management. In this work, we address this gap by first examining what constitutes a valuable reasoning experience, then introducing ExGRPO, a framework that strategically manages, selects and replays high-quality experiences. Experiments across multiple backbones and benchmarks show that ExGRPO consistently improves performance and stabilizes training where on-policy RLVR fails, demonstrating the importance of experience-aware optimization. We discuss limitations in Appendix A, and future work includes extending our method to multi-modal reasoning and agentic reinforcement learning.
@@
This research follows general code of ethics. We have taken care to ensure that the datasets, methodologies, and model used in our experiments are ethically sound. Our work is primarily theoretical and empirical, focusing on optimizing reasoning models. As such, it does not involve sensitive personal data, or content likely to introduce societal bias, harm, or discrimination.
@@
We provide the experimental setups and hyperparameters in Section 5.1, with further details in Appendix E. The code and weights of the ExGRPO series models are publicly available on GitHub and Hugging Face.
@@ Appendix A Limitations
While our work demonstrates the significant benefits of principled experience management in RLVR, we acknowledge several limitations. First, since our research scope focused on RLVR, the evaluation tasks are verifiable problems, i.e., mathematical and general reasoning benchmarks where trajectories can be more easily verified as correct or incorrect. The applicability of our correctness-based bucketing strategy to more open-ended tasks (e.g., creative writing), where rewards are often subjective and dense, remains an open question. Second, our framework’s definition of a valuable experience is based on heuristics that, while powerful, might be incomplete. ExGRPO might neglect the learning potential of “valuable failures” where some incorrect paths that contain useful training signals (Zhu et al., 2025). This focus on exploitation could risk premature convergence in some scenarios. Finally, while ExGRPO is built upon a relative policy optimization objective, its interaction with other families of RL algorithms has not been explored.
@@ D.1 Recap and Formulation
We will sometimes write $w(\cdot)$ or $W(\cdot)$ for brevity. For the replayed item $o^{*}$, ExGRPO uses a bounded policy-shaping transform $f(w)\,{=}\,\frac{w}{w+\beta}$ with $\beta>0$ instead of hard clipping; this is analogous to truncated/saturated importance ratios in ACER and V-trace, known to control variance at small bias (Wang et al., 2017; Espeholt et al., 2018). We assume bounded rewards, absolute continuity of policies (mutual support overlap), and an auxiliary KL control as in Schulman et al. (2015; 2017).

### v2

@@ 1 Introduction
Experience replay (Lin, 1992) is a widely adopted technique in RL to address this issue and improve sample efficiency. As stated in Silver & Sutton (2025), AI is at the cusp of a new period in which experience will become the dominant medium of improvement. This idea, often referred to as experience-based RL, leverages the intuition that previously collected interactions (i.e., state-action-reward tuples) contain valuable information for learning, helping models stabilize training and accelerate convergence (Mnih et al., 2013). A non-trivial challenge lies in how to exploit past experiences based on their differing “values” (Schaul et al., 2016) and manage the replay process according to customized learning schedules (Sujit et al., 2023). However, efficient experience replay mechanisms remain largely underexplored in RLVR for building large reasoning models (LRMs;  Plaat et al. 2024). Given the vast quantity of experiences collected during rollouts, a fundamental question remains: How can the reasoning model effectively exploit its own stream of experience to maximize learning toward scaling RL compute for LRMs?
@@ 2 Related Work
Reinforcement Learning with Verifiable Rewards (RLVR; Lambert et al. 2024) frames the language model as a policy that generates reasoning trajectories for verifiable tasks, with rewards provided by rule-based (Guo et al., 2025; Yu et al., 2025) or model-based verifiers (Su et al., 2025; Ma et al., 2025b; Chen et al., 2025a). Most RLVR methods adopt on-policy optimization (Schulman et al., 2017; Shao et al., 2024), which ensures stability but incurs high computational cost. Recent work has explored off-policy techniques (Kallus & Uehara, 2020; Meng et al., 2023) that incorporate historical or external data. Some approaches mix expert demonstrations with on-policy updates, e.g., off-policy policy gradients (Yan et al., 2025), SFT loss (Zhang et al., 2025b; Ma et al., 2025a), or knowledge distillation (Xu et al., 2025); others develop direct off-policy update rules for improved sample efficiency (Cohen et al., 2025; Roux et al., 2025; Arnal et al., 2025). However, they overlook the impact of data quality within the experience replay buffer, a factor that remains underexplored in RLVR and is the central focus of our work.
@@
Experience replay (Lin, 1992) is a classical RL technique to improve sample efficiency and stabilize training, later extended with prioritized replay (Schaul et al., 2016; Sujit et al., 2023) to emphasize informative transitions, enabling breakthroughs in control tasks (Mnih et al., 2013; 2015; Lillicrap et al., 2016). For LRMs, recent studies show that replaying successful trajectories accelerates convergence and improves reasoning capabilities. ReMix (Liang et al., 2025) leverages off-policy experience replay to improve training efficiency; RePO adopts online replay (Li et al., 2025), collecting early on-policy rollouts and revisiting them asynchronously; RLEP (Zhang et al., 2025a) reuses trajectories from a well-trained policy; and RRL (Dou et al., 2025) dynamically revisits promising early states; ARPO (Lu et al., 2025) extends this idea to GUI agents, combining GRPO with a replay buffer to enhance stability and sample efficiency. Among them, the last three methods overlook importance weighting to correct distribution mismatch in off-policy updates. In this work, we argue that not all stored experiences are equally valuable in zero RLVR for LRMs. We systematically analyze properties that determine the value of past experiences and design methods to better leverage high-value experiences, thereby improving efficiency and performance.
@@ 3.2 Preliminary Study on Experience Data
Figure 1: Analysis of question difficulty and entropy in on-policy RLVR training: (a) Test performance of models trained on different question groups; (b) Entropy comparison between correct and incorrect trajectories; (c) Entropy distributions of logically correct trajectories across question groups;
@@
Entropy magnitude (Chen et al., 2025b) is a key signal in RLVR training. Prior work shows that aggressively minimizing entropy can cause collapse (Cui et al., 2025b), motivating a more balanced exploration–exploitation strategy in experience management. While high-entropy policies encourage exploration, they can yield low-quality trajectories, many successes are merely lucky hits from incorrect reasoning chains.
@@ 4.1 ExGRPO: Experience Management
The replay buffer $\mathcal{E}$ is partitioned into buckets by each question’s latest correctness rate $\mathrm{Acc}(q^{*})$, with $q^{*}$ as the indexing key. Prior work (Zhang & Zuo, 2025) reweights advantages based on correctness rates, whereas we leverage them to bucket experiences by difficulty. To prevent overfitting on trivial cases, we introduce a Retired Set: questions solved in all rollouts are removed from $\mathcal{E}$, ensuring optimization focuses on partially solved questions. Xiong et al. (2025) also filters prompts by online correctness. In contrast, our retired-set strategy excludes mastered experience items.
@@
(2) Trajectory Selection. For each sampled question, we select the trajectory with the lowest entropy under the current policy without relying on an external judge, i.e., $o^{*}\leftarrow\arg\min_{o_{i}\in\{o^{*}\}}H(o_{i};\pi_{\theta})$, thereby prioritizing trajectories which indicate better CoT quality as discussed in Section 3.2.
@@ 4.2 ExGRPO: Experiential Policy Optimization
The on-policy objective $\mathcal{J}_{\text{on}}(\theta)$ follows GRPO in Eq. 3: for each query $q\sim\mathcal{B}_{\text{on}}$, $K$ trajectories $\{o_{i}\}_{i=1}^{K}$ are sampled from policy $\pi_{\theta_{\text{old}}}$ to form an advantage group $\mathcal{G}_{q}$, with objective $\mathcal{J}_{\text{GRPO}}(q,\mathcal{G}_{q};\theta)$.
@@ 5.2 Main Results
5.3 Extension to RL with Continuous Rewards
Beyond the binary-reward RLVR setup, extending ExGRPO to standard RLHF with continuous preference-model rewards is a vital next step. In this setting, we replace rollout success rates with the within-query reward variance computed over multiple rollouts. Reward variance serves as a natural signal of difficulty, and trajectory entropy remains applicable as a measure of reasoning validity.
For each query $q$ with $K$ rollouts, we compute the reward standard deviation $\sigma_{q}$. Since $\sigma_{q}$ has no intrinsic scale, we map it to a discrete difficulty bucket using relative, batch-wise normalization. Given all variances $\{\sigma_{1},\ldots,\sigma_{M}\}$ within the current batch, we apply a linear min-max normalization $\tilde{\sigma}_{q}=\frac{\sigma_{q}-\sigma_{\min}}{\max(\sigma_{\max}-\sigma_{\min},\varepsilon)},$ where $\varepsilon$ prevents degeneracy when the batch variance collapses. We then discretize $\tilde{\sigma}_{q}\in[0,1]$ into $\{0,\ldots,K_{\mathrm{rollout}}\}$ via $\operatorname{round}\!\left(\tilde{\sigma}_{q}\cdot K_{\mathrm{rollout}}\right).$ Each query is then assigned to its corresponding bucket and updated using the ExGRPO mechanism originally developed for the binary-correctness setting.
Table 2: Preliminary results extending ExGRPO to RL with continuous rewards.
Method ArenaHardV2 IFEval IFBench Avg.
GRPO 15.1 26.1 18.7 20.0
ExGRPO 16.4 27.9 20.7 21.7
Under this formulation, buckets still preserve the nature of discrete difficulty levels. As a result, ExGRPO’s sampling mechanism, which was previously centered around medium success rates, now preferentially selects queries with moderate reward variance, i.e., those that are neither overly stable nor excessively noisy. These correspond to intermediate-difficulty queries where learning gains are often most substantial.
To validate feasibility, we conducted a study on 7k WildChat-IF samples (Zhao et al., 2025; Bhaskar et al., 2025) with on-policy RL training (2 epochs), using Qwen2.5-7B-Base as the base model and Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) as the reward model. Results on chat and instruction-following benchmarks show a improvement of ExGRPO extension.
@@ 6 Analysis and Discussion
To understand the effectiveness of ExGRPO, we investigate the following research questions: 1) How does experience replay influence training dynamics, especially for different models? 2) What is the impact of experience selection on overcoming performance bottlenecks? 3) How do the components of ExGRPO, e.g., experience management, contribute to overall performance?
@@
6.1 Training Dynamics
Figure 5: Dynamics of experience replay buffer and retried set.
@@ 6.1 Training Dynamics
Figure 6: Dynamics of experience under different data conditions.
@@
The efficiency of experience utilization, rather than its sheer volume, is the primary driver of model performance. Figure 6 illustrates the dynamics of the experience buffer and the retired set under different data conditions. When trained on the same amount of data, the absence of question selection mechanisms (e.g., ExGRPO vs. w/o Ques. Sel.) leads to growth of the retired set and a corresponding reduction in buffer size (primarily due to repeated exposure to easy questions). As a result, experience replay becomes inefficient, with high consumption of samples but low return in terms of performance gains. We further examine the scaling effects of the experience replay ratio, $\rho$. A high ratio ($\rho=75\%$) leads to over-exploitation, stifling exploration and degrading performance below the on-policy baseline. Conversely, a low ratio ($\rho=25\%$) prioritizes fresh on-policy learning; however, while it enlarges the experience buffer, the retired set grows only marginally. This indicates that accumulating more experiences does not guarantee improved capabilities. Ultimately, effective replay relies on balancing exploration with efficient sample reuse rather than simply maximizing buffer size, with optimal performance achieved at a balanced $\rho=50\%$ (see results in Appendix Table 6).
@@ 6.2 Ablation Studies
We conduct an ablation study to assess the contributions of core components of ExGRPO. Figure 7 shows performance dynamics of validation set (see Section E.1). For both question and trajectory selection, our analytically derived heuristics consistently outperform random selection, while random selection itself offers marginal gains over the on-policy baseline. Policy shaping mitigates this by ensuring exploitation does not hinder exploration. It is also noteworthy that applying shaping alone in on-policy RLVR yields no improvement (Yan et al., 2025), confirming that ExGRPO’s gains cannot be attributed to naive shaping. The above results and additional ablations on test set (see Section E.4) provide empirical support for our design of ExGRPO.
Sensitivity to difficulty sampling strategies.
Our method employs a probabilistic sampling scheme based on a Gaussian distribution $p\propto\mathcal{N}(\mathrm{Acc}(q);\,\mu=0.5,\sigma)$. To assess the sensitivity of ExGRPO to the width of this distribution, we vary $\sigma\in\{0.5,1.5\}$. A smaller $\sigma$ corresponds to a narrower medium-difficulty band, whereas a larger $\sigma$ approaches a near-uniform sampling distribution. We also evaluate a Hard-Biased variant by shifting the mean to $\mu=0.25$, which emphasizes harder samples (25% accuracy) while reducing the chance of revisiting easier ones (75% accuracy). As shown in Table 3, the asymmetric sampling strategies yield inferior performance compared with the symmetric medium-difficulty focus ($\mu=0.5$), confirming that balanced uncertainty regions provide the most effective signal.
Table 3: Sensitivity analysis of the Gaussian sampler with varying sampling focus and bias. Bold indicates the best results across settings for each column.
Sampling Strategy In-Distribution Performance Out-of-Distribution Performance
AIME24 AIME25 AMC MATH Minerva Olympiad Avg. ARC-c GPQA MMLU-P Avg.
On-Policy 24.9 15.5 59.2 84.8 38.2 49.3 45.3 82.6 37.4 49.2 56.4
ExGRPO, ($\mu{=}0.5,\sigma{=}1$) 31.6 18.7 66.3 87.4 36.0 50.1 48.4 84.7 37.4 52.9 58.3
Wider Focus ($\sigma{=}1.5$) 22.0 19.2 62.3 85.4 38.2 49.9 46.2 81.7 44.4 48.6 58.2
Narrower Focus ($\sigma{=}0.5$) 24.9 17.4 64.9 86.6 34.2 49.0 46.2 72.9 40.4 51.1 54.8
Hard-Biased ($\mu{=}0.25$) 22.4 19.6 61.9 86.0 38.2 49.8 46.3 83.6 41.4 49.4 58.2
@@ 7 Conclusion
While experiential RL offers a promising approach to mitigate sample inefficiency in RLVR for large reasoning models, this area still lacks systematic exploration of experience value and management. In this work, we address this gap by first examining what constitutes a valuable reasoning experience, then introducing ExGRPO, a framework that strategically manages, selects and replays high-quality experiences. Experiments across multiple backbones and benchmarks show that ExGRPO consistently improves performance and stabilizes training where on-policy RLVR fails.
@@
This research follows the ICLR Code of Ethics. We have taken care to ensure that the datasets, methodologies, and model used in our experiments are ethically sound. Our work is primarily theoretical and empirical, focusing on optimizing reasoning models. As such, it does not involve sensitive personal data, or content likely to introduce societal bias, harm, or discrimination.
@@
We provide the experimental setups and hyperparameters in Section 5.1, with further details in Appendix E. The code and model weights are open-sourced.
Acknowledgments
We extend our gratitude to all the reviewers for their valuable feedback and suggestions. This work was supported by the Shanghai Artificial Intelligence Laboratory. This work was supported by the Science and Technology Development Fund of Macau SAR (Grant Nos. FDCT/0007/2024/AKP, EF2024-00185-FST), the UM and UMDF (Grant Nos. MYRG-GRG2024-00165-FST-UMDF, MYRG-GRG2025-00236-FST, SHMDF-AI/2026/001), and the National Natural Science Foundation of China (Grant No. 62266013).
@@ Appendix A Limitations
While our work demonstrates the significant benefits of principled experience management in RLVR, we acknowledge several limitations. First, since our research scope focused on RLVR, the evaluation tasks are verifiable problems, i.e., mathematical and general reasoning benchmarks where trajectories can be more easily verified as correct or incorrect. The applicability of our correctness-based bucketing strategy to more open-ended tasks (e.g., creative writing), where rewards are often subjective and dense, remains an open question. Second, our framework’s definition of a valuable experience is based on heuristics that, while powerful, might be incomplete. ExGRPO might neglect the learning potential of “valuable failures” where some incorrect paths that contain useful training signals (Zhu et al., 2025). This focus on exploitation could risk premature convergence in some scenarios. Finally, while ExGRPO is built upon a relative policy optimization objective, its interaction with other families of RL algorithms has not been explored. Future work includes extending our method to multi-modal reasoning and agentic reinforcement learning.
@@ D.1 Recap and Formulation
We will sometimes write $w(\cdot)$ or $W(\cdot)$ for brevity. For the replayed item $o^{*}$, ExGRPO uses a bounded policy-shaping transform $f(w)\,{=}\,\frac{w}{w+\beta}$ with $\beta>0$ instead of hard clipping; this is analogous to truncated/saturated importance ratios in ACER and V-trace, known to control variance at small bias (Wang et al., 2016; Espeholt et al., 2018). We assume bounded rewards, absolute continuity of policies (mutual support overlap), and an auxiliary KL control as in Schulman et al. (2015; 2017).